# An investigation of air pollution-induced temperature sensitivity and susceptibility to heat-related hospitalization in the Medicare population

### v1

@@ Abstract
Background: With rising temperatures and an aging US population, understanding how to prevent heat-related illness among older Americans will be an increasingly critical objective. Despite biological plausibility, no study to date has investigated how exposure to fine particulate matter air pollution (PM2.5) may contribute to risk of heat-related hospitalization.
Methods: We identified Medicare fee-for-service beneficiaries who experienced a heat-related hospitalization between 2008 and 2016. Using a case-crossover design and fitting Bayesian conditional logistic regression models, we characterized the association between heat-related hospitalization and temperature and PM2.5 exposures. We estimated the relative excess risk due to interaction (RERI) to quantify the additive-scale interaction of simultaneous exposure to heat and PM2.5.
Results: We observed 112,969 heat-related hospitalizations. Fixing PM2.5 at the case day median, increasing temperature from its case day median to the 95th percentile was associated with an odds ratio of 1.045 (95% CI: 1.026, 1.063). Fixing temperature at the case day median and increasing PM2.5 from its median to the 95th percentile was associated with an odds ratio of 1.014 (95% CI: 0.993, 1.037). We estimated the RERI associated with simultaneous median-to-95th percentile increases in temperature and PM2.5 to be 0.032 (0.007, 0.057).
Conclusion: Using nationwide Medicare claims and a self-matched study design, we found evidence supporting synergism between temperature and PM2.5 exposures on the risk of heat-related hospitalization.
Keywords: Heat-related illness; Temperature; Air pollution; Exposure synergism; Case-crossover; Medicare
@@ 1 Introduction
Heat is a leading cause of weather-related illness and death, and its impact is expected to intensify with climate change (Ebi et al.,, 2021). Previous research suggests that each year, thousands of deaths in the United States are directly or indirectly attributable to heat exposure (Weinberger et al.,, 2020). Extreme heat events have been associated with increased hospitalizations for conditions that may be triggered or exacerbated by heat, such as respiratory (Anderson et al.,, 2013), renal (Fletcher et al.,, 2012), and cardiovascular (Sun et al.,, 2018) diseases, as well as conditions that occur as a direct result of heat exposure (Layton et al.,, 2020), which are broadly referred to as heat-related illness. Heat-related illness ranges in severity from mild conditions such as heat cramps and syncope to severe, life-threatening conditions like heat stroke. Many of these outcomes are preventable through proactive measures such as avoiding direct sun, maintaining hydration, or temporarily relocating to public cooling centers—actions often prompted by heat alerts (Hopp et al.,, 2018). However, despite these preventive efforts, heat-related emergency department visits in the United States reached record levels during the summer of 2023 (Vaidyanathan,, 2024).
@@
Older adults are particularly vulnerable to heat due to impaired thermoregulation, social isolation, a higher prevalence of comorbidities, and the use of medications that increase heat sensitivity (Khalaj et al.,, 2010; Flynn et al.,, 2005; Layton et al.,, 2020). As the US population ages and extreme heat events become more frequent and intense (Pachauri et al.,, 2014), it will become increasingly essential to understand and implement strategies to protect older Americans from heat-related illness.
@@
Previous studies have found synergistic effects between acute exposure to heat and particulate matter (PM) air pollution on mortality (Stafoggia et al.,, 2008; Li et al.,, 2017; Kioumourtzoglou et al.,, 2016) and on certain cause-specific hospitalizations (Yitshak-Sade et al.,, 2018). However, despite biological plausibility, no study has yet investigated the role of PM in heat-related hospitalizations. Fine particulate matter (PM2.5) is small enough to penetrate the lungs and bloodstream, affecting several organ systems. Evidence from epidemiological and toxicological studies show that acute exposure to PM2.5 can impair vascular function, trigger oxidative stress and inflammation, and increase blood pressure (Pryor et al.,, 2022; Basith et al.,, 2022). These physiological responses may hinder the body’s ability to thermoregulate in hot conditions, potentially increasing the risk of heat-related illness when PM2.5 levels are high. Understanding the impact of PM2.5 on heat-related illness is critical for improving heat alert systems and guiding preparedness efforts.
@@
To address this gap, we investigate the independent and combined effects of temperature and PM2.5 on heat-related hospitalizations among older Medicare fee-for-service (FFS) beneficiaries. Specifically, our analysis focuses on inpatient hospitalizations for heat-related illness during the warm season (June–September) from 2008 to 2016. We employ a case-crossover design and a Bayesian analytic approach to achieve two main objectives: 1) reduce bias from unmeasured, time-invariant confounders (Basu et al.,, 2005; Maclure and Mittleman,, 2000), and 2) provide interpretable inferences using posterior distributions generated by Markov chain Monte Carlo (MCMC) sampling. We model nonlinear associations between the exposures and outcomes, and we assess synergy between heat and PM2.5 by estimating the relative excess risk due to interaction (RERI), an additive-scale interaction quantity.
@@ 2.1 Study population and case definition
We assessed records from a 50% random sample of Medicare FFS beneficiaries aged 65 and older who resided in the contiguous United States and were continuously enrolled in Parts A, B, and D for at least 12 months between 2008 and 2016. These records include data on beneficiaries’ sex, age, race/ethnicity, Medicaid eligibility status, ZIP code of residence, inpatient admission and discharge dates, and date of death, if applicable. We identified inpatient hospitalizations with a primary or secondary diagnosis of heat-related illness, based on International Classification of Diseases (ICD; 9th and 10th revisions) diagnosis codes. Detailed ICD-9 and -10 codes are provided in the supplemental materials (Table S1). We refer to the composite of these indications as heat-related hospitalization, or HRH. Our analysis is limited to individuals who experienced an HRH during the warm season (June–September), and we included only each beneficiary’s first HRH during the study period, excluding any subsequent events.
@@ 2.2 Exposure assessment
Daily maximum temperature data were obtained from the Parameter-elevation Relationships on Independent Slopes (PRISM) model (PRISM Climate Group,, 2023; Daly et al.,, 2008), which estimates temperature across the contiguous United States using surface station data and geographic factors such as elevation, coastal proximity, and atmospheric conditions (Daly et al.,, 2008). Daily temperature estimates are available at a 4 kilometer (km) resolution. For each ZIP code, we assigned the temperature estimate from the PRISM grid cell whose centroid was nearest to the ZIP code’s population weighted centroid.
Daily PM2.5 concentrations were derived from a validated ensemble model that combines ground monitoring data, chemical transport models, satellite observations, and land-use variables (Di et al.,, 2019, 2025). These estimates, available at a 1 km 1 km resolution, were aggregated to the ZIP code level by averaging grid cell values whose centroids fell within each ZIP code.
For each individual and date, exposure was defined as: 1) the daily maximum temperature on that date and 2) the three-day average PM2.5 concentration, calculated as the mean of the current and two preceding days. We used an extended exposure window for PM2.5 to account for the delayed health effects that might be incurred by air pollution exposures, which has been observed in previous studies (Wang et al.,, 2018; Fernández-Niño et al.,, 2018).
@@ 2.3 Statistical analysis
We used a case-crossover design to evaluate the effects of acute heat and PM2.5 exposures among individuals who experienced an HRH, with each individual serving as their own control. The case day was defined as the admission date of an individual’s first HRH during the study period. Control days were assigned bidirectionally within the same calendar month and matched to the same day of the week as the case day, with up to four matched control days per case selected before or after the hospitalization date. To eliminate potentially undue influence from extreme PM2.5 concentration outliers, we calculated the 95th percentile of three-day PM2.5 exposures across all case and control days and excluded observations exceeding this threshold (Di et al.,, 2017; Nethery et al.,, 2023).
@@
We implemented the Bayesian analogue to conditional logistic regression using the brms package (Bürkner,, 2017) to estimate the associations of temperature and PM2.5 with HRH. The primary model incorporated natural cubic splines for temperature and PM2.5, along with a linear interaction term. As a sensitivity analysis, we also applied a tensor product spline approach to enable a more complex, nonlinear interaction between the exposures. Detailed descriptions of both approaches are provided in the supplementary materials.
@@
We estimated the odds of HRH under various temperature and PM2.5 scenarios, reporting posterior means and 95% Bayesian credible intervals (CIs) derived from the posterior distribution samples of the MCMC. While the measures of association from conditional logistic regression models are typically evaluated on a multiplicative scale, interpreting multiplicative interactions can be challenging to interpret in a public health context (Rothman,, 2012). Instead, we assess the interaction between heat and PM2.5 on an additive scale by estimating the relative excess risk due to interaction (RERI) (Knol et al.,, 2007; VanderWeele and Knol,, 2014). The RERI quantifies the increased risk associated with simultaneous exposure to heat and PM2.5 beyond the sum of their individual effects. Because HRH is a rare outcome, the odds ratio-based RERI estimate should closely approximate the conventional RERI, which is a function of relative risks (VanderWeele and Knol,, 2014).
@@
To estimate the RERI for two continuous exposures, we select the median and 95th percentile values from the distributions of temperature and three-day average PM2.5 concentrations on case days as contrast levels. This approach to assessing synergism between exposures has been applied in similar epidemiological studies (Fayyad et al.,, 2024; Josey et al.,, 2023). Additional methodological details are provided in the supplemental materials. Finally, we also report the interactions between temperature and PM2.5 on a multiplicative scale as recommended by VanderWeele and Knol, (2014) and Vandenbroucke et al., (2007), estimating this measure using posterior means and 95% Bayesian CIs from the linear interaction coefficient.
@@
All analyses were completed in R version 4.4.0 (R Core Team,, 2021); code to reproduce these analyses is available at https://github.com/laurenmock/temp_pm25_heat-hosp.
@@ 3.1 Main Results
Among 29,345,820 Medicare beneficiaries in our sample, we observed at least one heat-related hospitalization for 112,969 unique individuals, who served as cases in our case-crossover study. The characteristics of the individuals who experienced an HRH are detailed in Table 1. Most of the hospitalized individuals were women (68.0%) and white (82.3%), with 22.9% eligible for Medicaid. The mean age at the time of hospitalization was 81.7 years. On case days, the mean maximum daily temperature was 84.6F with the 5th and 95th percentile exposures at 68.7F and 98.4F, respectively. Prior to trimming, the case day distribution of three-day PM2.5 concentrations was heavily right skewed with exposures as high as 120.4 . Case and control days with three-day PM2.5 exposures above 18.4 , the 95th percentile of the distribution, were trimmed; this resulted in a mean PM2.5 concentration of 9.4 on case days. PM2.5 exposure levels were notably high among Black beneficiaries and low among American Indian/Alaska Native beneficiaries (Table 1). The full distributions of temperature and PM2.5 case day exposures are displayed in Figures S1 and S2 in the supplementary materials.
| Case day exposures | |||
| Mean (SD) | |||
| Daily max. | Three-day | ||
| N (%) | temperature (F) | PM2.5 () | |
| Total | 112,969 | 84.6 (9.0) | 9.4 (3.7) |
| Sex | |||
| Male | 35,930 (31.8%) | 84.9 (9.1) | 9.4 (3.7) |
| Female | 77,039 (68.2%) | 84.5 (9.0) | 9.5 (3.7) |
| Age category | |||
| 65-74 | 19,165 (17%) | 85.1 (9.1) | 9.6 (3.7) |
| 75-84 | 45,868 (40.6%) | 84.8 (9.0) | 9.4 (3.7) |
| 85+ | 47,936 (42.4%) | 84.3 (9.0) | 9.4 (3.7) |
| Race/ethnicity | |||
| White | 92,975 (82.3%) | 84.4 (9.0) | 9.3 (3.7) |
| Black/African-American | 13,123 (11.6%) | 85.9 (8.2) | 10.3 (3.5) |
| Asian/Pacific Islander | 2,006 (1.8%) | 82.7 (9.4) | 9.8 (3.6) |
| Hispanic | 3,178 (2.8%) | 87.8 (9.6) | 9.3 (3.4) |
| American Indian/Alaska Native | 650 (0.6%) | 86.6 (10.3) | 7.6 (3.8) |
| Other | 1,037 (0.9%) | 84.4 (9.5) | 9.6 (3.6) |
| Medicaid eligibility | |||
| Ineligible | 87,059 (77.1%) | 84.5 (9.0) | 9.3 (3.7) |
| Eligible | 25,910 (22.9%) | 85.1 (9.1) | 9.9 (3.6) |
@@
Temperature exposure on case days was highest in the South and Southwest regions of the US (Figure 1, left panel), while PM2.5 exposure concentrations were highest in the South, Midwest, and Southern California (Figure 1, middle panel). The rates of HRH per 1,000 Medicare FFS beneficiaries at risk varied substantially across the United States, with the highest rates in the Gulf Coast and Great Plains regions (Figure 1, right panel).
@@
Figure 2 visualizes the independent nonlinear effects of heat and PM2.5 on heat-related hospitalization; we hold PM2.5 at its case day median to examine the effect of heat, and hold temperature at its case day median to examine the effect of PM2.5. Maximum daily temperature exhibited a strong positive association with odds of HRH (Figure 2, left panel). Specifically, when we fix PM2.5 concentrations at 8.9 , increasing the temperature from 85.3F (median) to 98.4F (95th percentile) yields an odds ratio (95% CI) of 1.05 (1.03, 1.06) (Figure 3 and Table S2 in the supplementary materials). Although a positive association was observed between three-day PM2.5 and HRH, credible intervals were wide and covered one across the range of PM2.5 levels (Figure 2, right panel). When we fix temperature at 85.3F, an increase in PM2.5 concentrations from 8.9 (median) to 16.1 (95th percentile) corresponds to an odds ratio (95% CI) of 1.01 (0.99, 1.04) (Figure 3 and Table S2 in the supplementary materials).
@@
We observed an additive interaction between temperature and PM2.5. A simultaneous increase in both temperature and PM2.5 from their median values to their respective 95th percentile quantities was associated with an odds ratio (95% CI) of 1.09 (1.06, 1.12) (Figure 3 and Table S2 in the supplementary materials). Using these odds ratios to approximate the RERI (95% CI), we found that the approximate risk increase was 0.03 (0.01, 0.06) (Figure 3 and Table S2 in the supplementary materials), indicating synergism between the two exposures resulting in excess heat-related hospitalizations. We visualized synergism between these two exposures by comparing the odds of HRH at median exposure levels to the odds of HRH across a grid of exposure values (Figure 4). Temperature is strongly associated with HRH, but on hot days, the odds of HRH also increases with higher concentrations of PM2.5. On cooler days, however, we see a small protective effect of PM2.5.
@@
Although we were primarily interested in assessing the additive interaction with the RERI, we also estimated the interaction on a multiplicative scale. The estimated odds ratio associated with the linear interaction term was 1.00031 (95% CI: 1.00006, 1.00056). Both the additive and multiplicative interaction estimates were positive with credible intervals that did not cover the null value.
3.2 Sensitivity analysis

### v2

# Excess risk of heat-related hospitalization associated with temperature and PM2.5 among older adults

@@ Abstract
Background: With rising temperatures and an aging population, understanding how to prevent heat-related illness among older adults will be increasingly crucial. Despite biological plausibility, no study to date has investigated how exposure to fine particulate matter air pollution (PM2.5) may contribute to the risk of hospitalization with a diagnosis code indicating heat-related illness, referred to as heat-related hospitalization. This study aims to fill this gap by investigating the independent and combined effects of temperature and PM2.5 exposures on heat-related hospitalization risk.
Methods: We identified Medicare fee-for-service beneficiaries in the contiguous United States who experienced a heat-related hospitalization between 2008 and 2016. Using a case-crossover design and fitting Bayesian conditional logistic regression models, we characterized the associations of temperature and PM2.5 exposures with heat-related hospitalization. We then estimated the relative excess risk due to interaction to quantify the additive-scale interaction of simultaneous exposure to heat and PM2.5.
Results: We observed 112,969 heat-related hospitalizations among 29,345,820 Medicare beneficiaries in the study sample. Fixing PM2.5 at the case day median, the odds ratio for increasing temperature from its case day median to the 95th percentile was 1.05 (95% CI: 1.03, 1.06). Fixing temperature at the case day median, the odds ratio for increasing PM2.5 from its median to the 95th percentile was 1.01 (95% CI: 0.99, 1.04). The risk due to interaction for simultaneous median-to-95th percentile increases in temperature and PM2.5 was 0.03 (95% CI: 0.01, 0.06).
Conclusions: Our study is the first to observe synergism between temperature and PM2.5 exposures associated with the risk of heat-related hospitalization. These findings highlight the importance of considering air pollution in effective public health and clinical interventions to prevent heat-related illness.
@@ 1 Introduction
Heat is a leading cause of weather-related illness and death, and its impact is expected to intensify with climate change. 1 Previous research suggests that each year, thousands of deaths in the United States are directly or indirectly attributable to heat exposure. 2 Extreme heat events have been associated with increased hospitalizations for conditions that may be triggered or exacerbated by heat, such as respiratory, 3 renal, 4 and cardiovascular 5 diseases, as well as conditions that occur as a direct result of heat exposure, 6 which are broadly referred to as heat-related illness. Heat-related illness ranges in severity from mild conditions such as heat cramps and syncope to severe, life-threatening conditions like heat stroke. Many of these outcomes are preventable through proactive measures such as avoiding direct sun, maintaining hydration, or temporarily relocating to public cooling centers—actions often prompted by heat alerts. 7; 8 However, despite these preventive efforts, heat-related emergency department visits in the United States reached record levels during the summer of 2023. 9
@@
Older adults are particularly vulnerable to heat due to impaired thermoregulation, social isolation, a higher prevalence of comorbidities, and the use of medications that increase heat sensitivity. 6; 8; 10 As the US population ages and extreme heat events become more frequent and intense, 11 it will become increasingly essential to understand and implement strategies to protect older Americans from heat-related illness.
@@
Previous studies have identified synergistic effects between acute exposure to heat and particulate matter (PM) air pollution on mortality 12; 13; 14 and on certain cause-specific hospitalizations. 15 These investigations of synergism are especially critical since daily temperature and PM2.5 levels are often correlated across much of the United States, even after accounting for seasonal and long-term trends, leading to frequent simultaneous high exposures. 16 Despite biological plausibility, no study has yet investigated the role of PM in heat-related hospitalizations (HRH). PM2.5 is small enough to penetrate the lungs and bloodstream, affecting several organ systems. Evidence from epidemiological and toxicological studies shows that acute exposure to PM2.5 can impair vascular function, trigger oxidative stress and inflammation, and increase blood pressure. 17; 18 These physiological responses may hinder the body’s ability to thermoregulate in hot conditions, potentially increasing the risk of heat-related illness when PM2.5 levels are high. Understanding the impact of PM2.5 on heat-related illness is critical for improving heat alert systems and guiding preparedness efforts.
@@
To address this gap, we investigate the independent and combined effects of temperature and PM2.5 on HRHs, defined as hospitalizations due to heat-related illness, among older Medicare fee-for-service (FFS) beneficiaries. Specifically, our analysis focuses on inpatient hospitalizations for heat-related illness during the warm season (June–September) from 2008 to 2016. We employ a case-crossover design and a Bayesian analytic approach to: 1) reduce bias from unmeasured, time-invariant confounders, 19; 20 and 2) provide interpretable inferences using posterior distributions generated by Markov chain Monte Carlo sampling. We model nonlinear associations between the exposures and outcomes, and we assess synergy between heat and PM2.5 by estimating the relative excess risk due to interaction (RERI), an additive-scale interaction quantity.
@@ 2.1 Study population and case definition
We assessed records from a 50% random sample of Medicare FFS beneficiaries aged 65 and older who resided in the contiguous United States and were continuously enrolled in Parts A, B, and D for at least 12 months between 2008 and 2016. These records include data on beneficiaries’ sex, age, race/ethnicity, Medicaid eligibility status, ZIP code of residence, inpatient admission and discharge dates, and date of death, if applicable. We identified inpatient hospitalizations with a primary or secondary diagnosis of heat-related illness, based on International Classification of Diseases (9th and 10th revisions) diagnosis codes; detailed International Classification of Diseases-9 and -10 codes are provided in Table S1 of Supplementary Material. In the United States, the majority of unscheduled inpatient hospitalizations follow an emergency room visit, although some admissions originate with physician referrals from outpatient settings. 21 We refer to the composite of these indications as HRH. Our analysis is limited to individuals who experienced an HRH during the warm season (June–September), and we included only each beneficiary’s first HRH during the study period, excluding any subsequent events.
@@ 2.2 Exposure assessment
Daily maximum temperature data were obtained from the Parameter-elevation Relationships on Independent Slopes Model, 22; 23 which estimates temperature across the contiguous United States using surface station data and geographic factors such as elevation, coastal proximity, and atmospheric conditions. Daily temperature estimates are available at a 4-km resolution. For each ZIP code, we assigned the temperature estimate from the Parameter-elevation Relationships on Independent Slopes Model grid cell whose centroid was nearest to the ZIP code’s population-weighted centroid. Daily PM2.5 concentrations were derived from a validated ensemble model that integrates ground monitoring data, chemical transport models, satellite observations, and land-use variables. 24; 25 These estimates, available at a 1-km resolution, were aggregated to the ZIP code level by averaging values from grid cells whose centroids fell within each ZIP code.
For each individual and date, exposure was defined as: 1) the daily maximum temperature in the individual’s residential ZIP code on that date, and 2) the three-day average PM2.5 concentration in the individual’s residential ZIP code, calculated as the mean concentration from the case day and the two preceding days. We used an extended exposure window for PM2.5 to account for the delayed health effects that might be incurred by air pollution exposures, which has been observed previously. 26; 27 As some previous studies have identified lagged effects of heat on hospitalizations, 28; 29 we conducted a sensitivity analysis using three-day exposure windows for both temperature and PM2.5.
@@ 2.3 Statistical analysis
We used a case-crossover design to evaluate the effects of acute heat and PM2.5 exposures among individuals who experienced an HRH, with each individual serving as their own control. The case day was defined as the admission date of an individual’s first HRH during the study period. Control days were assigned bidirectionally within the same calendar month and matched to the same day of the week as the case day, with up to four matched control days per case selected before or after the hospitalization date. To eliminate potentially undue influence from extreme PM2.5 concentration outliers, we calculated the 95th percentile of three-day PM2.5 exposures across all case and control days and trimmed, or excluded, observations exceeding this threshold. 30; 31 We also conducted a sensitivity analysis using the 99th percentile as the threshold for trimming.
@@
We implemented the Bayesian analogue to conditional logistic regression using the brms package 32 to estimate the associations of temperature and PM2.5 with HRH. The primary model comprised three terms: a natural cubic spline for temperature, a natural cubic spline for PM2.5, and a linear interaction term. As a sensitivity analysis, we also applied a tensor product spline approach to enable a more complex, nonlinear interaction between the exposures. Detailed descriptions of both approaches are provided in the Supplementary Material.
@@
We estimated the odds of HRH under various temperature and PM2.5 scenarios, reporting posterior means and 95% Bayesian credible intervals (CIs) derived from the posterior distribution samples of the Markov chain Monte Carlo. While the measures of association from conditional logistic regression models are typically evaluated on a multiplicative scale, multiplicative interactions can be challenging to interpret in a public health context. 33 Instead, we assess the interaction between heat and PM2.5 on an additive scale by estimating the relative excess risk due to interaction (RERI). 34; 35 The RERI quantifies the increased risk associated with simultaneous exposure to heat and PM2.5 beyond the sum of their individual effects. Because HRH is a rare outcome, the odds ratio-based RERI estimate should closely approximate the conventional RERI, which is a function of relative risks. 35
@@
To estimate the RERI for two continuous exposures, we select the median and 95th percentile values from the distributions of temperature and three-day average PM2.5 concentrations on case days as contrast levels. This approach to assessing synergism between exposures has been applied in similar epidemiological studies. 36; 37 Additional methodological details are provided in the Supplementary Material. Finally, we also report the interactions between temperature and PM2.5 on a multiplicative scale as recommended by VanderWeele and Knol 35 as well as Vandenbroucke et al, 38 estimating this measure using posterior means and 95% Bayesian CIs from the linear interaction coefficient.
@@
All analyses were completed in R (version 4.4.0; Vienna, Austria); 39 code to reproduce these analyses is available at https://github.com/laurenmock/temp_pm25_heat-hosp.
Our study aims and data stewardship plan were approved by the Rutgers University Institutional Review Board (Pro20170001685). Medicare patient data are protected by our data usage agreement with the Centers for Medicare and Medicaid, and were stored and analyzed with minimal risk of breach of data confidentiality. Use of Medicare data does not require informed consent from individual beneficiaries.
@@ 3.1 Main analysis
Among 29,345,820 Medicare beneficiaries in our sample, we observed at least one HRH among 112,969 unique individuals. Each individual’s first HRH served as a case in our case-crossover study. Heat-related illness was the primary diagnosis for 60,053 hospitalizations (53.2%) and the secondary diagnosis for the remaining hospitalizations. The mean length of stay in the hospital was 3.5 (standard deviation: 4.5) days. The characteristics of the individuals who experienced an HRH are detailed in Table 1. Most of the hospitalized individuals were women (68.2%) and white (82.3%), with 22.9% eligible for Medicaid. The mean age at the time of hospitalization was 81.7 years. On case days and control days, the mean maximum daily temperatures were 29.2C and 29.1C, respectively. Before trimming, the case day distribution of three-day PM2.5 concentrations was heavily right skewed, with exposures as high as 120.4 . Case and control days with three-day PM2.5 exposures above 18.4 , the 95th percentile of the distribution, were trimmed; this resulted in a mean PM2.5 concentration of 9.4 on case days and 9.3 on control days. The full distributions of temperature and PM2.5 case day exposures are displayed in Figures S1 and S2 in the Supplementary Material.
| N (%) | |
| Total | 112,969 |
| Sex | |
| Male | 35,930 (31.8%) |
| Female | 77,039 (68.2%) |
| Age category | |
| 65-74 | 19,165 (17.0%) |
| 75-84 | 45,868 (40.6%) |
| 85+ | 47,936 (42.4%) |
| Race/ethnicity | |
| White | 92,975 (82.3%) |
| Black/African-American | 13,123 (11.6%) |
| Hispanic | 3,178 (2.8%) |
| Asian/Pacific Islander | 2,006 (1.8%) |
| American Indian/Alaska Native | 650 (0.6%) |
| Other | 1,037 (0.9%) |
| Medicaid eligibility | |
| Ineligible | 87,059 (77.1%) |
| Eligible | 25,910 (22.9%) |
| Mean (SD) | |
| Daily maximum temperature (C) | |
| Case days | 29.2 (5.0) |
| Control days | 29.1 (5.1) |
| Three-day PM2.5 () | |
| Case days | 9.4 (3.7) |
| Control days | 9.3 (3.6) |
SD, standard deviation.
@@
Temperature exposure on case days was highest in the South and Southwest regions of the United States (Figure 1, left panel), while PM2.5 exposure concentrations were highest in the South, Midwest, and Southern California (Figure 1, middle panel). The rates of HRH per 1000 Medicare FFS beneficiaries at risk varied substantially across the United States, with the highest rates in the Central United States (Figure 1, right panel).
@@
Figure 2 shows the independent effects of heat and PM2.5 on HRH, with PM2.5 held at its case day median to examine the effect of heat, and temperature held at its case day median to examine the effect of PM2.5. Maximum daily temperature exhibited a strong positive association with odds of HRH (Figure 2, left panel). Specifically, when PM2.5 concentrations are fixed at 8.9 , increasing the temperature from 29.6C (median) to 36.9C (95th percentile) yields an odds ratio of 1.05 (95% CI: 1.03, 1.06) (Figure 3 and Table S2 in the Supplementary Material). Although a positive association was observed between three-day PM2.5 and HRH, credible intervals were wide and covered one across the range of PM2.5 levels (Figure 2, right panel). When temperature was fixed at 29.6C, an increase in PM2.5 concentrations from 8.9 (median) to 16.1 (95th percentile) yielded an odds ratio of 1.01 (95% CI: 0.99, 1.04) (Figure 3 and Table S2 in the Supplementary Material).
@@
Furthermore, we observed an additive interaction between temperature and PM2.5. A simultaneous increase in both temperature and PM2.5 from their median values to their respective 95th percentile quantities yielded an odds ratio of 1.09 (95% CI: 1.06, 1.12) (Figure 3 and Table S2 in the Supplementary Material). Using these odds ratios to approximate the RERI, we found that the approximate risk increase was 0.03 (95% CI: 0.01, 0.06) (Figure 3 and Table S2 in the Supplementary Material), indicating synergism between the two exposures resulting in excess HRHs. We visualized synergism between these two exposures by comparing the odds of HRH at median exposure levels to the odds of HRH across a grid of exposure values (Figure 4). Temperature was strongly associated with HRH, but on hot days, the odds of HRH also increased with higher concentrations of PM2.5. On cooler days, however, we see a small protective effect of PM2.5. Since temperature and PM2.5 were often correlated (as shown in Figure S2), we have limited data to assess the effect of elevated PM2.5 at lower temperatures during the warm season.
@@
Although we were primarily interested in assessing the additive interaction with the RERI, we also estimated the interaction on a multiplicative scale. The estimated odds ratio for the linear interaction term was 1.0003 (95% CI: 1.0001, 1.0006). Both the additive and multiplicative interaction estimates were positive, with credible intervals that did not cover the null value.
3.2 Sensitivity analyses
To investigate the interaction between temperature and PM2.5 further and assess the robustness of our findings, we fit three sensitivity analyses. Note that we used the median and 95th percentile contrasts defined in the main analysis across all sensitivity analyses to allow for direct comparisons across models.